What are the freezing point and boiling point of a solution that contains 5 g of glycerol (C3H8O3) and 20 ml of water?


1
Expert's answer
2021-03-09T06:04:30-0500

n(glycerol) = m/M = 5/60 = 0.0833 mol

m(H2O) = 20g = 0.02 kg

ΔT = E*n/m = 0.52*0.0833/0.02 = 2.1658°C

Boiling point is 102.1658°C

ΔT = K*n/m = 1.86*0.0833/0.02 = 7.7469°C

Freezing point is -7.7469°C